Postage rates. The maximum size envelope that can be mailed with a large envelope rate is 3 inches longer than it is wide. The are is 180in^2. Find the length and width.
Answer by scott8148(6628)

let x=width; so x+3=length

area = x(x+3) = 180

x^2 + 3x - 180 = 0

factoring ___ (x + 15)(x - 12) = 0

x +15 = 0 ___ x = -15 ___ negative value not realistic

x - 12 = 0 ___ x = 12 ___ this is the width

substituting ___ length = x + 3 = 15